A common mistake when solving equations is the following:
The equation: 2(x- 2) = x+ 3
First step in solving: 2x- 2= x + 3
Write a clear explanation of what error has been made. What could be done to
avoid this error?

Answer by vertciel(183)

2(x - 2) = x + 3
2(X - 2) means 2 multiplied by everything inside the bracket. You cannot just eliminate the bracket without multiplying.
So 2(x - 2) = 2x - 4.
I'll let you figure out how you can avoid this; it's easy, just use your common sense.